# muography; non-destructive testing; machine learning; image processing; upsampling; semantic segmentation; simulation; cosmic rays

### v1

@@ Abstract
The civil engineering industry faces a critical need for innovative non-destructive evaluation methods, particularly for ageing critical infrastructure, such as bridges, where current techniques fall short. Muography, a non-invasive imaging technique, constructs three-dimensional density maps by detecting the interactions of naturally occurring cosmic-ray muons within the scanned volume. Cosmic-ray muons offer both deep penetration capabilities due to their high momenta and inherent safety due to their natural source. However, the technology’s reliance on this natural source results in a constrained muon flux, leading to prolonged acquisition times, noisy reconstructions and challenges in image interpretation. To address these limitations, we developed a two-model deep learning approach. First, we employed a conditional Wasserstein generative adversarial network with gradient penalty (cWGAN-GP) to perform predictive upsampling of undersampled muography images. Using the structural similarity index measure (SSIM), 1-day sampled images were able to match the perceptual qualities of a 21-day image, while the peak signal-to-noise ratio (PSNR) indicated a noise improvement to that of 31 days’ worth of sampling. A second cWGAN-GP model, trained for semantic segmentation, was developed to quantitatively assess the upsampling model’s impact on each of the features within the concrete samples. This model was able to achieve segmentation of rebar grids and tendon ducts embedded in the concrete, with respective Dice-Sørensen accuracy coefficients of 0.8174 and 0.8663. This model also revealed an unexpected capability to mitigate — and in some cases entirely remove — z-plane smearing artifacts caused by the muography’s inherent inverse imaging problem. Both models were trained on a comprehensive dataset generated through Geant4 Monte-Carlo simulations, designed to reflect realistic civil infrastructure scenarios. Our results demonstrate significant improvements in both acquisition speed and image quality, marking a substantial step toward making muography more practical for reinforced concrete infrastructure monitoring applications.
@@ 1 Introduction
It has been widely established that the growing quantity of reinforced-concrete-based infrastructure nearing the end of its original intended service life poses a significant challenge. Current reflection-based near-surface detection techniques — ground-penetrating radar (GPR) and ultrasonic echo measurements — are limited in scenarios with high concrete thickness and crowded imaging volumes. While X-ray planar tomography can provide high-resolution imaging at depth, its use is heavily restricted by radiation protection regulations meaning that it goes largely unused. Consequently, there are no established non-destructive testing (NDT) technologies capable of inspecting deep into concrete structures to determine placement of steel rebar reinforcement grids and tendon ducts, as well as identifying defects, such as honeycombing, tendon duct strand corrosion and air voiding.
@@
Muon scattering tomography, an NDT technique which utilises naturally occurring high-energy (GeV) cosmic ray muons as its imaging source, has shown promise in addressing this gap. Successfully deployed in applications such as nuclear waste characterisation Mahon et al. (2019), cargo scanning Barnes et al. (2023), and industrial pipe inspection Martinez Ruiz Del Arbol et al. (2022), muography has more recently been explored as a tool for civil engineering applications, including the inspection of bridges and other built infrastructure. A comparative study by Niederleithinger et al. Niederleithinger et al. (2021) evaluated GPR, ultrasound, X-rays and muography through the imaging of a 600 kg reference block of reinforced concrete, demonstrating that muography could identify the same features as benchmark X-rays while outperforming the resolution of GPR and ultrasound. However, challenges remain: muography requires days to weeks to produce an accurate tomograph caused by the low muon flux (1 muon $\mathrm{cm}^{-2}\mathrm{min}^{-1}$), compared to hours for GPR and ultrasound. Additionally, detector acceptance and limited vertical resolution caused by the inverse imaging problem also result in noisy images and shadowing artefacts from objects above and below the image plane.
@@
While computationally complex statistical methods Schultz et al. (2007) and pattern recognition techniques, including density-based clustering Giammanco et al. (2024) and support vector machines Das et al. (2022), have been proposed to enhance muographic imaging, they often fail to address the core issues of long imaging times, smearing effects and noisy outputs. Adjacent fields, such as medical imaging, have overcome some of these issues by using Image-to-Image (I2I) machine learning for image reconstruction and enhancement for Computed Tomography (CT), Magnetic Resonance Imaging (MRI) and Positron Emission Tomography (PET) Armanious et al. (2020); Zhou et al. (2021). While I2I applications are used to reduce imaging times and doses in hospital imaging, machine learning is also deployed as a detection tool using segmentation to aid medical professionals in detection of cancers, organs and polyps Ma et al. (2024).
@@
We can apply similar machine learning techniques to our muography images, however unlike medical imaging there is a lack of sufficient real-world data, which is required for training models. To address this, the Monte Carlo modelling software Geant4 Agostinelli et al. (2003); Allison et al. (2006, 2016), widely used in muography for proof-of-concept and feasibility studies, was employed. Geant4 allows for the simulation of the detectors, concrete blocks and muons, bypassing the need for costly experimental setups and their associated lengthy integration times. Using Geant4 along with the Ecomug cosmic muon event generator Pagano et al. (2021), a realistic dataset of high- and low-sampled muographic images of concrete interiors were generated for model training. Preliminary results demonstrate that a conditional Wasserstein generative adversarial network with gradient penalty (cWGAN-GP) model can effectively reduce noise and upsample features such as rebar grids and tendon ducts embedded in concrete. To analyse the perceptual improvements from upsampling, a second model was trained using the same architecture for a semantic segmentation task to enable pixelwise object detection; thus providing a quantified breakdown of the enhancements on each concrete feature as a result of upsampling process. The models significantly reduced noise, smoothed image edges and mitigated shadow artefacts caused by low vertical resolution. This work highlights the potential of machine learning in the post-processing of muographic images, addressing key challenges in the field. By enhancing image quality and reducing imaging artifacts, these advancements make muography a more viable and attractive technology for investment and industry adoption.
@@ 1.1 Muographic Imaging for Built Infrastructure
Muon scattering tomography and muon absorption radiography — collectively referred to as muography — are next-generation imaging techniques that leverage naturally occurring, high-energy cosmic-ray muons. These offer a novel NDT solution that overcomes the poor depth resolution and safety concerns of existing methods. Absorption measurements utilise the attenuation of cosmic-ray muons as they pass through matter, with lower detection rates indicating higher density regions of the imaged volume. This technique only requires a single tracker (consisting of a minimum of two detection planes) adjacent to or behind the target volume; thus it is often used for the imaging of large scale structures such as volcanoes Tanaka et al. (2014) and pyramids Morishima et al. (2017). Scattering tomography (hereafter referred to as muography) examines the multiple Coulomb scattering of muons as they interact with atomic nuclei while traversing matter. By analysing the variance in the muon scattering angle, the radiation length of the material along the muon’s path can be estimated. The true distribution of muon multiple scattering angles follows a Cauchy-Lorentz profile, but a Gaussian approximation can effectively describe 98% of the distribution, leading to the Rossi formula Rossi and Greisen (1941) for the standard deviation of muon multiple scattering angles:
$\sigma_{\theta}\approx\frac{15MeV}{\beta cp}\sqrt{\frac{L}{X_{0}}},$ (1)
where $\beta\mathrm{c}$ is velocity ($\beta\mathrm{=}1$), $\mathrm{L}$ is the length of matter traversed by the muon and $\mathrm{X_{0}}$ is the radiation length of the material. The measurement of the momentum $p$ (MeV/c) of a detected muon is dependent on the muography detector, with cost-effective detectors electing to use the average muon momentum ( $3-4\,\mathrm{GeV/c}$) over measurement. Simplifying Equation 1 for a non momentum-measurement, we find that the variance of the muon scattering angles is inversely proportional to the radiation length along the muons path:
@@
Muography thus aims to calculate the scattering angles of muons as they traverse a volume of interest (VoI). This measurement requires a minimum of four detector planes (in pairs of two) on either side of the VoI, measuring the ingoing and outgoing muon vectors.
Muography is inherently an inverse imaging problem, as due to multiple scattering, it is impossible to infer the exact path taken by the muon through the VoI. Various reconstruction algorithms have been developed over the years — detailed reviews of which can be found at Yang et al. (2019); Ughade et al. (2023) — however the Point of Closest Approach (PoCA) algorithm Schultz (2003) remains the most widely used for its simplicity and computational efficiency. Like most reconstruction algorithms, PoCA assumes one scattering event per muon which is calculated as the angle between the incoming and outgoing muon vectors, with the scattering point placed halfway along the vector describing the distance of closest approach between these vectors. Due to low muon flux and the ignorance of multiple scattering by these algorithms, accurate tomographs require many muon events. Therefore integration times, of the order of days and weeks, are required in order to gather enough muons events to statistically and sufficiently resolve objects embedded within thick structures such as concrete.
@@ 1.2 Convolutional Machine Learning Techniques for Image Processing
Supervised machine learning has emerged as a powerful data-driven tool capable of constructing models that are trained to generate outputs $\hat{y}$, which approximate ground truths $y=f(x)$. Machine learning is often only applied when the problem is too complex for conventional methods to solve, where it is able to identify obscure patterns and structures inherent to the dataset.
@@
Model training consists of two key steps: the forward pass and the backward pass. During the forward pass, inputs $x$ are passed through the parameters of the model to produce predictions $\hat{y}$. The accuracy of these predictions – measured by their similarity to the ground truth $y$ – is quantified by a differentiable loss function $\mathcal{L}(y,\hat{y})$, which evaluates the error between the predicted and actual outputs. This is followed by the backward pass, where the gradient of the loss function with respect to each model parameter is calculated. These gradients are then used by an optimisation function to update model parameters accordingly. This process is repeated thousands of times during training, with model parameters being iteratively adjusted to minimise the global loss, to achieve $\hat{y}\approx y$.
@@
A trained model can, at best, only generalise to the dataset on which it was trained; therefore it is crucial that the dataset accurately reflects the inputs the fully-trained model is expected to encounter. A well-trained model relies on a dataset that is representative of unseen data; as well as being diverse, expansive and sufficiently large to prevent underfitting, the model fails to capture necessary features resulting in poor performance on new data. Conversely, overfitting is often caused by an overly complex model which learns the specifics of the training dataset too well thus failing to generalise to unseen data. However, we can penalise overfit models so that they are more generalised — called regularisation — using common techniques such as non-linear activation functions, batch normalisation and dropout. Model performance is assessed on an unseen dataset (validation dataset) for continued assessment of model generalisation over the course of training, which is used to assess and guide hyperparameter tuning. Once training is complete, the fully-trained model is assessed on the test dataset — independent of the training and validation sets — to compare performance of different models and architectures.
@@
There are two primary mechanisms that can be used within image processing: convolution and attention. Convolutions have long been a cornerstone of image processing, even before the rise of modern machine learning, enabling the extraction of localised features from image data. These methods rely on pre-defined kernels to generate feature maps from input images. For instance, Sobel filters Sobel (2014) can be commonly used to detect vertical and horizontal edges in images. Convolutional neural networks (CNNs) build upon this concept by using learnable kernels, consisting of weight parameters which are tuned during training to extract intricate features and patterns from the data. CNNs typically consist of cascading convolutional layers, each containing multiple filters which act on the feature maps produced by the previous layers. This iterative process enables CNNs to learn increasingly abstract and detailed features, making them highly effective for a wide range of image processing tasks, including denoising, segmentation, classification, object detection, super-resolution, data fusion and image synthesis. However, CNNs have a fundamental limitation: their small receptive field. The information accessible to any one component is restricted to the spatial extent of the square convolutional kernel, typically sized between 3$\times$3 and 7$\times$7 pixels. Consequently, CNNs can only process localised areas of the input image at a time. In contrast, attention mechanisms can capture a global context, enabling the model to account for long-range dependencies across an image. Despite their advantages, attention mechanisms are more complex than CNNs and, given their relative recency, are less tested and explored in the imaging domain. For this initial study on the application of machine learning to muography, the focus will remain on well-established CNN architectures due to their proven reliability and effectiveness in the image processing domain.
1.2.1 The Conditional Generative Adversarial Network (cGAN)
@@
An encoder-decoder network is a model methodology that is widely used for CNNs. The encoder aims to learn the features that should be extracted from the input image, where the decoder aims to build the feature maps back up to an output image, enhancing relevant features and reducing redundant information. However, this is a lossy process so information deemed redundant to the network is lost, resulting in a loss of finer detail.
@@
In order to address the loss of detail in encoder-decoder convolutional networks, a 2015 paper by Ronneberger et al. Ronneberger et al. (2015) proposed a new architecture called a U-Net. They solve the problem of information loss by introducing skip connections between the encoder and decoder layers by cropping feature maps from the encoder and concatenating them to complimentary-sized feature maps in the decoder. This allows for minor details to be preserved, producing a higher quality output image. This architecture is widely used in image-to-image translation tasks Siddique et al. (2021), due to its simplicity when compared to alternative designs.
@@
A conditional Generative Adversarial Network (cGAN) consists of a generator that produces outputs and a discriminator that moderates these outputs during training, resulting in more accurate results than using a standalone U-Net generator. Whereas cGANs are conditional on an input, GANs by contrast are unsupervised generative models that aim to create data from a source of random noise. A standard GAN comprises a generator, which transforms the noise into an output, and a discriminator, which evaluates whether the output is generated data or part of the real dataset. The generator’s goal is to produce data that is indistinguishable from the real data, while the discriminator strives to correctly identify whether the data is real or fake. This dynamic creates an adversarial process, with the generator attempting to ‘fool’ the discriminator and the discriminator learning to better distinguish real from fake. The back-and-forth continues during training until a Nash equilibrium is reached — a state in game theory where neither player (the generator or the discriminator) can improve their position without affecting the other. In this context, the Nash equilibrium is achieved when the generator produces outputs realistic enough that the discriminator can not reliably differentiate between real and fake data and is forced to guess. Despite their powerful capabilities, GANs are notoriously difficult to train, often suffering from issues such as mode collapse (where the generator produces limited variability) or imbalance between the generator and discriminator (where one dominates the training process). To mitigate these challenges, techniques like the Wasserstein GAN and gradient penalty are frequently employed, improving training stability and convergence.

### v2

@@ Abstract
The civil engineering industry faces a critical need for innovative non-destructive evaluation methods, particularly for ageing critical infrastructure, such as bridges, where current techniques fall short. Muography, a non-invasive imaging technique, constructs three-dimensional density maps by detecting the interactions of naturally occurring cosmic-ray muons within the scanned volume. Cosmic-ray muons offer both deep penetration capabilities due to their high momenta and inherent safety due to their natural source. However, the technology’s reliance on this natural source results in a constrained muon flux, leading to prolonged acquisition times, noisy reconstructions, and challenges in image interpretation. To address these limitations, we developed a two-model deep learning approach. First, we employed a conditional Wasserstein Generative Adversarial Network with Gradient Penalty (cWGAN-GP) to perform predictive upsampling of undersampled muography images. Using the Structural Similarity Index Measure (SSIM), 1-day sampled images were able to match the perceptual qualities of a 21-day image, while the Peak Signal-to-Noise Ratio (PSNR) indicated a noise improvement to that of 31 days worth of sampling. A second cWGAN-GP model, trained for semantic segmentation, was developed to quantitatively assess the upsampling model’s impact on each of the features within the concrete samples. This model was able to achieve segmentation of rebar grids and tendon ducts embedded in the concrete, with respective Dice–Sørensen accuracy coefficients of 0.8174 and 0.8663. This model also revealed an unexpected capability to mitigate—and in some cases entirely remove—z-plane smearing artifacts caused by the muography’s inherent inverse imaging problem. Both models were trained on a comprehensive dataset generated through Geant4 Monte Carlo simulations designed to reflect realistic civil infrastructure scenarios. Our results demonstrate significant improvements in both acquisition speed and image quality, marking a substantial step toward making muography more practical for reinforced concrete infrastructure monitoring applications.
@@ 1 Introduction
It has been widely established that the growing quantity of reinforced-concrete-based infrastructure nearing the end of its original intended service life poses a significant challenge. Current reflection-based near-surface detection techniques—Ground- Penetrating Radar (GPR) and ultrasonic echo measurements—are limited in scenarios with high concrete thickness and crowded imaging volumes. While X-ray planar tomography can provide high-resolution imaging at depth, its use is heavily restricted by radiation protection regulations, meaning that it goes largely unused. Consequently, there are no established Non-Destructive Testing (NDT) technologies capable of inspecting deep into concrete structures to determine the placement of steel rebar reinforcement grids and tendon ducts, as well as identifying defects, such as honeycombing, tendon duct strand corrosion, and air voiding.
@@
Muon scattering tomography, an NDT technique which utilises naturally occurring high-energy (GeV) cosmic-ray muons as its imaging source, has shown promise in addressing this gap. Successfully deployed in applications such as nuclear waste characterization Mahon et al. (2019), cargo scanning Barnes et al. (2023), and industrial pipe inspection Martinez Ruiz Del Arbol et al. (2022), muon scattering techniques have been more recently explored as a tool for civil engineering applications, including the inspection of bridges and other built infrastructure such as historical buildings Bonomi et al. (2019). A comparative study by Niederleithinger et al. Niederleithinger et al. (2021) evaluated GPR, ultrasound, X-rays, and muography through the imaging of a 600 kg reference block of reinforced concrete, demonstrating that muography could identify the same features as benchmark X-rays while outperforming the resolution of GPR and ultrasound. However, challenges remain: muography requires days to weeks to produce an accurate tomograph caused by the low muon flux (1 muon $\mathrm{cm}^{-2}\,\mathrm{min}^{-1}$) compared to hours for GPR and ultrasound. Additionally, detector acceptance and limited vertical resolution caused by the inverse imaging problem also result in noisy images and shadowing artefacts from objects above and below the image plane.
@@
While computationally complex statistical methods Schultz et al. (2007) and pattern recognition techniques, including density-based clustering Giammanco et al. (2024) and support vector machines Das et al. (2022), have been proposed to enhance muographic imaging, they often fail to address the core issues of long imaging times, smearing effects, and noisy outputs. Adjacent fields, such as medical imaging, have overcome some of these issues by using Image-to-Image (I2I) machine learning for image reconstruction and enhancement for Computed Tomography (CT), Magnetic Resonance Imaging (MRI), and Positron Emission Tomography (PET) Armanious et al. (2020); Zhou et al. (2021). While I2I applications are used to reduce imaging times and doses in hospital imaging, machine learning is also deployed as a detection tool using segmentation to aid medical professionals in the detection of cancers, organs, and polyps Ma et al. (2024).
@@
We can apply similar machine learning techniques to our muography images; however, unlike medical imaging, there is a lack of sufficient real-world data, which are required for training models. To address this, the Monte Carlo modelling software Geant4 Agostinelli et al. (2003); Allison et al. (2006, 2016), widely used in muography for proof-of-concept and feasibility studies, was employed. Geant4 allows for the simulation of the detectors, concrete blocks, and muons, bypassing the need for costly experimental setups and their associated lengthy integration times. Using Geant4 along with the Ecomug cosmic muon event generator Pagano et al. (2021), a realistic dataset of high- and low-sampled muographic images of concrete interiors was generated for model training. Preliminary results demonstrate that a conditional Wasserstein Generative Adversarial Network with Gradient Penalty (cWGAN-GP) model can effectively reduce noise and upsample features such as rebar grids and tendon ducts embedded in concrete. To analyse the perceptual improvements from upsampling, a second model was trained using the same architecture for a semantic segmentation task to enable pixelwise object detection, thus providing a quantified breakdown of the enhancements on each concrete feature as a result of the upsampling process. The models significantly reduced noise, smoothed image edges, and mitigated shadow artefacts caused by low vertical resolution. This work highlights the potential of machine learning in the post-processing of muographic images, addressing key challenges in the field. By enhancing the image quality and reducing imaging artifacts, these advancements make muography a more viable and attractive technology for investment and industry adoption.
@@ 1.1 Muographic Imaging for Built Infrastructure
Muon scattering tomography and muon absorption radiography—collectively referred to as muography—are next-generation imaging techniques that leverage naturally occurring, high-energy cosmic-ray muons. These offer a novel NDT solution that overcomes the poor depth resolution and safety concerns of existing methods. Absorption measurements utilise the attenuation of cosmic-ray muons as they pass through matter, with lower detection rates indicating higher density regions of the imaged volume. This technique only requires a single tracker (consisting of a minimum of two detection planes) adjacent to or behind the target volume. Therefore, it is suited to a wide variety of tasks, having been demonstrated on built infrastructure—such as railway tunnels Thompson et al. (2020), dams Oláh et al. (2023), and reactor buildings Fujii et al. (2020)—as well as larger structures such as volcanoes Tanaka et al. (2014) and pyramids Morishima et al. (2017).
Scattering tomography (hereafter referred to as muography) examines the multiple Coulomb scattering of muons as they interact with atomic nuclei while traversing matter. This technique requires a minimum of four detector planes arranged as two pairs on either side of the volume of interest, which limits its practical applications, yet it remains preferred over absorption methods due to its superior spatial resolution and detail. By analyzing the variance in the muon scattering angle, the radiation length of the material along the muon’s path can be estimated. The true distribution of muon multiple scattering angles follows a Cauchy–Lorentz profile, but a Gaussian approximation can effectively describe 98% of the distribution, leading to the Rossi formula Rossi and Greisen (1941) for the standard deviation of muon multiple scattering angles:
$\sigma_{\theta}\approx\frac{15\,\mathrm{MeV}}{\beta cp}\sqrt{\frac{L}{X_{0}}},$ (1)
[…]
@@
Positional detector hits can be used to calculate ingoing and outgoing muon vectors as they traverse the VoI, allowing the scattering angle to be determined. Using the above relationships, material composition can then be inferred from the scattering angles.
Muography is inherently an inverse imaging problem, as due to multiple scattering, it is impossible to infer the exact path taken by the muon through the VoI. Various reconstruction algorithms have been developed over the years—detailed reviews of which can be found at Yang et al. (2019); Ughade et al. (2023)—however, the Point of Closest Approach (PoCA) algorithm Schultz (2003) remains the most widely used for its simplicity and computational efficiency. Like most reconstruction algorithms, the PoCA assumes one scattering event per muon, which is calculated as the angle between the incoming and outgoing muon vectors, with the scattering point placed halfway along the vector describing the distance of closest approach between these vectors. Due to low muon flux and the ignorance of multiple scattering by these algorithms, accurate tomographs require many muon events. Therefore, integration times, of the order of days and weeks, are required in order to gather enough muons events to statistically and sufficiently resolve objects embedded within thick structures such as concrete.
@@ 1.2 Convolutional Machine Learning Techniques for Image Processing
Supervised machine learning has emerged as a powerful data-driven tool capable of constructing models that are trained to generate outputs $\hat{y}$, which approximate ground truths defined as $y=f(x)$. Machine learning is often only applied when the problem is too complex for conventional methods to solve, where it is able to identify obscure patterns and structures inherent to the dataset.
@@
Model training consists of two key steps: the forward pass and the backward pass. During the forward pass, inputs $x$ are passed through the parameters of the model to produce predictions $\hat{y}$. The accuracy of these predictions—measured by their similarity to the ground truth $y$—is quantified by a differentiable loss function $\mathcal{L}(y,\hat{y})$, which evaluates the error between the predicted and actual outputs. This is followed by the backward pass, where the gradient of the loss function with respect to each model parameter is calculated. These gradients are then used by an optimization function to update model parameters accordingly. This process is repeated thousands of times during training, with model parameters being iteratively adjusted to minimise the global loss to achieve $\hat{y}\approx y$.
@@
A trained model can, at best, only generalise to the dataset on which it was trained; therefore, it is crucial that the dataset accurately reflects the inputs that the fully trained model is expected to encounter. A well-trained model relies on a dataset that is representative of unseen data; in addition to being diverse, expansive, and sufficiently large to prevent underfitting, the model fails to capture necessary features, resulting in poor performance on new data. Conversely, overfitting is often caused by an overly complex model which learns the specifics of the training dataset well, thus failing to generalise to unseen data. However, we can penalise overfit models so that they are more generalised—called regularisation—using common techniques such as non-linear activation functions, batch normalization, and dropout. Model performance is assessed on an unseen dataset (validation dataset) for continued assessment of the model generalization over the course of training, which is used to assess and guide hyperparameter tuning. Once training is complete, the fully trained model is assessed on the test dataset—independent of the training and validation sets—to compare the performance of different models and architectures.
@@
There are two primary mechanisms that can be used within image processing: convolution and attention. Convolutions have long been a cornerstone of image processing even before the rise of modern machine learning, enabling the extraction of localised features from image data. These methods rely on predefined kernels to generate feature maps from input images. For instance, Sobel filters Sobel (2014) can be commonly used to detect vertical and horizontal edges in images. Convolutional Neural Networks (CNNs) build upon this concept by using learnable kernels consisting of weight parameters which are tuned during training to extract intricate features and patterns from the data. CNNs typically consist of cascading convolutional layers, each containing multiple filters which act on the feature maps produced by the previous layers. This iterative process enables CNNs to learn increasingly abstract and detailed features, making them highly effective for a wide range of image processing tasks, including denoising, segmentation, classification, object detection, super-resolution, data fusion, and image synthesis. However, CNNs have a fundamental limitation: their small receptive field. The information accessible to any one component is restricted to the spatial extent of the square convolutional kernel, which is typically sized between 3 $\times$ 3 and 7 $\times$ 7 pixels. Consequently, CNNs can only process localised areas of the input image at a time. In contrast, attention mechanisms can capture a global context, enabling the model to account for long-range dependencies across an image. Despite their advantages, attention mechanisms are more complex than CNNs and, given their relative recency, are less tested and explored in the imaging domain. For this initial study on the application of machine learning to muography, the focus will remain on well-established CNN architectures due to their proven reliability and effectiveness in the image processing domain.
The Conditional Generative Adversarial Network (cGAN)
@@
An encoder–decoder network is a model methodology that is widely used for CNNs. The encoder aims to learn the features that should be extracted from the input image, where the decoder aims to build the feature maps back up to an output image, enhancing relevant features and reducing redundant information. However, this is a lossy process so information deemed redundant to the network is lost, resulting in a loss of finer detail.
@@
In order to address the loss of detail in encoder–decoder convolutional networks, a 2015 paper by Ronneberger et al. Ronneberger et al. (2015) proposed a new architecture called a U-Net. They solved the problem of information loss by introducing skip connections between the encoder and decoder layers by cropping feature maps from the encoder and concatenating them to complimentary-sized feature maps in the decoder. This allows for minor details to be preserved, producing a higher quality output image. This architecture is widely used in image-to-image translation tasks Siddique et al. (2021) due to its simplicity when compared to alternative designs.
@@
A conditional Generative Adversarial Network (cGAN) consists of a generator that produces outputs and a discriminator that moderates these outputs during training, resulting in more accurate results than using a standalone U-Net generator. Whereas cGANs are conditional on an input, GANs by contrast are unsupervised generative models that aim to create data from a source of random noise. A standard GAN comprises a generator, which transforms the noise into an output, and a discriminator, which evaluates whether the output has generated data or part of the real dataset. The generator’s goal is to produce data that are indistinguishable from the real data, while the discriminator strives to correctly identify whether the data are real or fake. This dynamic creates an adversarial process, with the generator attempting to ‘fool’ the discriminator and the discriminator learning to better distinguish real from fake data.
The back-and-forth continues during training until a Nash equilibrium is reached—a state in game theory where neither player (the generator or the discriminator) can improve their position without affecting the other. In this context, Nash equilibrium is achieved when the generator produces outputs that are realistic enough that the discriminator can not reliably differentiate between real and fake data and is forced to guess. Despite their powerful capabilities, GANs are notoriously difficult to train, often suffering from issues such as mode collapse (where the generator produces limited variability) or imbalance between the generator and discriminator (where one dominates the training process). To mitigate these challenges, techniques like the Wasserstein GAN and gradient penalty are frequently employed, improving training stability and convergence.